OpenStudy (anonymous):

state the gernal term. t4=35 and common difference is -12

OpenStudy (anonymous):

\(t_4=35\) \(t_3=47\) \(t_2=59\) \(t_1=62\) i think, because the common difference is 12

OpenStudy (anonymous):

so in general \(t_n=73-12n\) i suppose, although maybe this is not what you want

OpenStudy (anonymous):

answer says it should be tn=83-12n but i dont know how to get that

OpenStudy (anonymous):

cause i messed up sorry

OpenStudy (anonymous):

i took the fourth term and kept adding 12 to count down, but my addition was sloppy

OpenStudy (anonymous):

\(t_4=35\) \(t_3=35+12=47\) \(t_2=47+12=59\) \(t_1=59+12=71\) \[t_0=71+12=83\]

OpenStudy (anonymous):

subtract 12 each time, the \(n\)th term will look like \(t_n=83-12n\)

OpenStudy (anonymous):

thanks :)